# {cutout}00.3cm22cm1 empty    PointMamba: A Simple State Space Model for Point Cloud Analysis

### v2

@@ 5 Conclusion
In this study, we propose a state space model (SSM) for point cloud analysis, named PointMamba. One challenge with applying unidirectional SSM is that it cannot capture non-causal data like point clouds. By taking embedded point patches as input and applying a simple reordering strategy, we adapt point cloud data for the unidirectional data-dependent SSM, i.e., Mamba. We also attempted to pre-train PointMamba, further expanding the potential of our method. Experimental results show that PointMamba outperforms the transformer-based counterparts on different point cloud analysis datasets while significantly reducing parameters and FLOPs. As a newcomer to point cloud analysis, PointMamba is a promising option for constructing 3D vision foundation models, and we hope it can offer a new perspective for the field.

### v3

@@ 5 Conclusion
In this study, we propose a state space model (SSM) for point cloud analysis named PointMamba. One challenge with applying unidirectional SSM is that it cannot capture non-causal data like point clouds. By taking embedded point patches as input and applying a simple reordering strategy, we adapt point cloud data for the unidirectional data-dependent SSM, i.e., Mamba. We also attempted to pre-train PointMamba, further expanding the potential of our method. Experimental results show that PointMamba outperforms the transformer-based counterparts on different point cloud analysis datasets while significantly reducing parameters and FLOPs. As a newcomer to point cloud analysis, PointMamba is a promising option for constructing 3D vision foundation models, and we hope it can offer a new perspective for the field.